OpenStudy (anonymous):

use L'Hospital Rule. lim x->0+ (tan5x)^x

OpenStudy (fibonaccichick666):

What do you need to use Mr. L'hops?

OpenStudy (dumbcow):

\[(\tan 5x)^x = e^{\ln (\tan 5x)^x} = e^{x \ln(\tan 5x)}\] now apply the limit as x->0

OpenStudy (anonymous):

would the limit be 1?

OpenStudy (anonymous):

Thank you so much.

OpenStudy (dumbcow):

yw yes lim = 1

OpenStudy (fibonaccichick666):

but that's not using l'hopital?

OpenStudy (dumbcow):

is it required to use L'Hopitals rule?

OpenStudy (anonymous):

no, if a more elementary way can be used than...

OpenStudy (fibonaccichick666):

it said so

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

so, since we have to use L'hops, what do we need to use it?